If f(x)=
5x+3
----- ,then f(c+2)=
x+5
NOTE- 5x+3 is divided by x+5

Answer by stanbon(48522)

If f(x)=
5x+3
----- ,then f(c+2)=
x+5
===================
f(c+2) = [(c+2)+3] / [(c+2)+5]
f(c+2) = (c+5)/(c+5)
---------
Cheers,
Stan H.